Solve for x.
x^2+32=8x
I got (8+8i/2 , 8-8i/2)
Am I on the right track?
Thank you

Answer by stanbon(75887)

Solve for x.
x^2+32=8x
---
x^2-8x+32 = 0
---
x = [8 +- sqrt(64-128)]/2
---
x = [8 +- sqrt(-64)]/2
---
x = [8 +- 8i]/2
---
x = (4+4i) or x = (4-4i)
==========================
Cheers,
Stan H.
===============